# Integral cohomology rings of weighted Grassmann orbifolds and rigidity properties

### 11 Mar 2024 v2

@@ 2.2. $q$-CW complex structures of weighted Grassmann orbifolds
(2.11) $R(\lambda^{i})=\{\lambda^{j}:\lambda^{j}=(\lambda^{i}_{s},r)\lambda^{i}\text{ for }(\lambda^{i}_{s},r)\in{\rm{rev}}(\lambda^{i})\text{ and }s=1,2,\dots,k\}.$
Note that $|R(\lambda^{i})|=d(\lambda^{i})\leq i$ for all $i=0,1,\dots,m$. Consider $\mathbb{C}^{d(\lambda^{i})}\subset\mathbb{C}^{m+1}$ by
$\mathbb{C}^{d(\lambda^{i})}=\{(z_{0},z_{1},\dots,z_{m})~{}|~{}z_{j}=0\text{ if}\lambda^{j}\notin R(\lambda^{i})\}.$
Then, ${E}(\lambda^{i})$ is homeomorphic to $\mathbb{C}^{d(\lambda^{i})}$ by the following map
$g\colon{E}(\lambda^{i})\to\mathbb{C}^{d(\lambda^{i})}.$
The $j$-th entry of $g([a_{1}\wedge\dots\wedge a_{k}])$ is equal to $a_{sr}$ (the $r$-th entry of $a_{s}$) if $\lambda^{j}=(\lambda^{i}_{s},r)\lambda^{i}$ for some $1\leq s\leq k$, $1\leq r\leq n$ and $r\notin\{\lambda^{i}_{1},\dots,\lambda^{i}_{k}\}$. This defines both $g$ and $g^{-1}$ completely. Note that this is a bijective correspondence, and both $g$ and $g^{-1}$ are continuous.
Let $\widetilde{E}(\lambda^{i})=\pi^{-1}(E(\lambda^{i}))$. Then
$\widetilde{E}(\lambda^{i})=\{(z_{0},z_{1},\dots,z_{m})\in Pl(k,n)~{}|~{}z_{i}\neq 0,z_{j}=0\text{ for }j>i\}.$
Note that $\widetilde{E}(\lambda^{i})$ is homeomorphic to ${E}(\lambda^{i})\times\mathbb{C}^{*}$ via the map
The inverse map is defined by
Now $\widetilde{E}(\lambda^{i})\subset Pl(k,n)$ and there is a there is a restricted $\bf b$-action of $\mathbb{C}^{*}$ on $\widetilde{E}(\lambda^{i})$. Define a ${\bf{b}}$-action on ${E}(\lambda^{i})\times\mathbb{C}^{*}$ by
$t([z],s)=([tz],t^{b_{i}}s).$
Then $f$ is $\mathbb{C}^{*}$-equivariant with respect to the above ${\bf{b}}$-actions.
Note that $g$ is also $\mathbb{C}^{*}$-equivariant map, where the action of $\mathbb{C}^{*}$ on $\mathbb{C}^{d(\lambda^{i})}$ is restriction of the ${\bf{b}}$-action on $\mathbb{C}^{m+1}$. Then
$(g,\mbox{id})\circ f:\widetilde{E}(\lambda^{i})\to\mathbb{C}^{d(\lambda^{i})}\times\mathbb{C}^{*}$
is an $\mathbb{C}^{*}$-equivariant homeomorphism. Let $G(b_{i})$ be the group of $b_{i}$-th roots of unity. The finite subgroup $G(b_{i})$ of $\mathbb{C}^{*}$ acts on the second factor trivially. Thus
$\pi_{\bf{b}}(\widetilde{E}(\lambda^{i}))\cong\frac{\mathbb{C}^{d(\lambda^{i})}\times\mathbb{C}^{*}}{{\bf b}\text{-action}}\cong\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}.$
Thus the $q$-cell decomposition of $\mbox{Gr}_{\bf b}(k,n)$ is given by $\mbox{Gr}_{\bf{b}}(k,n)=\sqcup_{i=0}^{m}\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}.$
For each $j\in\{0,1,\dots,m\}$, define $X_{j}=\sqcup_{i=0}^{j}\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}.$ The building sequence corresponding to the $q$-CW complex structure is
Next, we recall lens complex from [13]. Let $(c_{0},c_{1},\dots,c_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ and $G(c_{m})$ be the $c_{m}$-th roots of unity.
Definition 2.11.
[13] Consider an action of $G(c_{m})$ on $S^{2m-1}$ defined by
$t(z_{0},z_{1},\dots,z_{m-1})=(t^{c_{0}}z_{0},t^{c_{1}}z_{1},\dots,t^{c_{m-1}}z_{m-1}),$
where $t\in G(c_{m})$. The quotient space
$L^{\prime}(c_{m};(c_{0},c_{1},\dots,c_{m-1})):=\dfrac{S^{2m-1}}{G(c_{m})}$
is called the lens complex. If $c_{0},\dots,c_{m-1}$ are prime to $c_{m}$ then it is a usual generalized lens space.
Now consider a Plücker weight vector ${\bf b}=(b_{0},b_{1},\dots,b_{m})$. For every $i=1,2,\dots,m$ define a weight vector
(2.13) ${\bf{b}}^{(i)}:=\{(b_{0},b_{1},\dots,b_{i-1})\in(\mathbb{Z}_{\geq 1})^{d(\lambda^{i})}\text{ such that }b_{j}=0\text{ if }\lambda^{j}\notin R(\lambda^{i})\}.$
We denote the following lens complex by $L^{\prime}(b_{i};{\bf{b}}^{(i)})$
$L^{\prime}(b_{i};{\bf{b}}^{(i)}):=\{[z_{0},z_{1},\dots,z_{i-1}]\in L^{\prime}(b_{i};(b_{0},b_{1},\ldots,b_{i-1}))~{}|~{}z_{j}=0\text{ if }\lambda^{j}\notin R(\lambda^{i})\}$
@@
Consider $\mathbb{C}^{d(\lambda^{i})}=\{(z_{0},z_{1},\dots,z_{m})~{}|~{}z_{j}=0\text{ if}\lambda^{j}\notin R(\lambda^{i})\}\subset\mathbb{C}^{m}$ and denote the unit circle on $\mathbb{C}^{d(\lambda^{i})}$ by $S^{2d(\lambda^{i})-1}=\{z\in\mathbb{C}^{d(\lambda^{i})}~{}|~{}|z|=1\}$. Then
$L^{\prime}(b_{i};{\bf{b}}^{(i)})=\dfrac{S^{2d(\lambda^{i})-1}}{G(b_{i})}.$
Using the $q$-cell decomposition $X^{i}=X^{i-1}\sqcup\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}$. Now the result follows from the fact that $\frac{\mathbb{C}^{d(\lambda^{i})}}{G(b_{i})}$ is an unbounded cone over $L^{\prime}(b_{i};{\bf{b}}^{(i)})$. ∎
@@ 3. Classification of weighted Grassmann orbifolds upto certain homeomorphism
In this section, we introduce the Plücker permutation that induces a permutation on Plücker coordinates up to some sign. Then we prove that a Plücker permutation induces an equivariant homeomorphism between two weighted Grassmann orbifolds. We also show that two weighted Grassmann orbifolds are homeomorphic if their Plücker weight vectors are differed by scalar multiplication. Conversely, we prove that if two weighted Grassmann orbifolds are homeomorphic with a small assumption, their weight vectors differ by scalar multiplication and a permutation. This provides an equivariant classification of a weighted Grassmann orbifold in terms of Plücker weight vectors.
@@
A permutation $\sigma$ on $\{0,1,2,\dots,m\}$ is said to be Plücker permutation if there exist $t_{\sigma}\in\{1,-1\}^{m+1}$ such that $t_{\sigma}{\cdot}\sigma z\in Pl(k,n)$ for every $z\in Pl(k,n)$.
Let $\sigma$ and $\eta$ be two Plücker permutations on $\{0,1,2,\dots,m\}$. Then $\sigma\circ\eta$ is also a Plücker permutation on $\{0,1,2,\dots,m\}$. Note that, $\sigma^{-1}=\sigma^{c}$ for some positive integer $c$. Thus $\sigma^{-1}$ is also a Plücker permutation on $\{0,1,2,\dots,m\}$. Hence the Plücker permutations on $\{0,1,\dots,m\}$ form a subgroup of the permutation group on $\{0,1,\dots,m\}$.
@@
  1. (1)
    Let $\lambda^{i}=(\lambda_{1}^{i},\ldots,\lambda_{k}^{i})$ be the $i$-th Schubert symbol described in (2.2), for $i=0,\ldots,m$ and $\phi\in S_{n}$. Define $\phi\lambda^{i}:=(\phi(\lambda_{i_{1}}^{i}),\ldots,\phi(\lambda_{i_{k}}^{i}))$, where $i_{1},\dots,i_{k}\in\{1,\dots,k\}$ such that $1\leq\phi(\lambda_{i_{1}}^{i})<\phi(\lambda_{i_{2}}^{i})<\cdots<\phi(\lambda_{i_{k}}^{i})\leq n$. Then $\phi\lambda^{i}\in\{\lambda^{0},\dots,\lambda^{m}\}$. Thus $\phi\in S_{n}$ induces a permutation $\sigma$ on $\{0,1,\dots,m\}$ defined by $\phi\lambda^{i}=\lambda^{\sigma(i)}$. Then $\sigma$ is a Plücker permutation on $\{0,1,\dots,m\}$ follows from (2.4). Thus there exists an injective group homomorphism from $S_{n}$ to the group of the Plücker permutations on $\{0,1,\dots,m\}$ if $m+1={n\choose k}$.
  2. (2)
    The converse is not always true. For a particular case $n=4,k=2$, the total number of Plücker permutations are $48$. Among them, a total of $24$ permutations forms $S_{4}$.
@@
Let ${\bf{b}}=(b_{0},\dots,b_{m})$ be a Plücker weight vector. For every $(z_{0},\dots,z_{m})\in Pl(k,n)$ we have $(t^{b_{0}}z_{0},\dots,t^{b_{m}}z_{m})\in Pl(k,n)$ fot every $t\in\mathbb{C}^{*}$. If $\sigma$ is a Plücker permutation on $\{0,1,\dots,m\}$ then there exist $t_{\sigma}=(t_{0},\dots,t_{m})\in\{1,-1\}^{m+1}$ such that $(t_{0}z_{\sigma(0)},\dots,t_{m}z_{\sigma(m)})$ is in $Pl(k,n)$. Moreover, $t_{\sigma}{\cdot}\sigma(t^{b_{0}}z_{0},\dots,t^{b_{m}}z_{m})$ is also a Plücker coordinate. Thus $(t_{0}t^{b_{\sigma(0)}}z_{\sigma(0)},\dots,t_{m}t^{b_{\sigma(m)}}z_{\sigma(m)})\in Pl(k,n)$. Then using Remark 2.2, it follows that $\sigma{\bf{b}}=(b_{\sigma(0)},\dots,b_{\sigma(m)})$ is a Plücker weight vector. ∎
Now we can consider the weighted Grassmann orbifolds $\mbox{Gr}_{\sigma{\bf{b}}}(k,n)$ for a Plücker permutations $\sigma$. Plücker permutations play an important role in characterizing the weighted Grassmann orbifold up to a weakly equivariant homeomorphism. Consider the (*$(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\bf{b}}(k,n)$ defined by
(3.1) $(t_{1},t_{2},\dots,t_{n})\sum_{i=0}^{m}[a_{i}e_{\lambda^{i}}]=\sum_{i=0}^{m}[t_{\lambda^{i}}a_{i}e_{\lambda^{i}}]$
@@
Theorem 3.4.
[…]
@@
The homeomorphism $f_{\sigma}$ becomes weakly equivariant with respect to the permuted (*$(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\sigma{\bf b}}(k,n)$ defined by
(3.2) $(t_{1},t_{2},\dots,t_{n})\sum_{i=0}^{m}[a_{i}e_{\lambda^{i}}]=\sum_{i=0}^{m}[t_{\lambda^{\sigma(i)}}a_{i}e_{\lambda^{i}}]$
Now we discuss how the Plücker permutation $\sigma$ affects the $q$-CW complex structure on $\mbox{Gr}_{\bf{b}}(k,n)$. The permutation of the coordinates in $\mathbb{C}^{m+1}$ induces another decomposition of $Pl(k,n)$.
(3.3) $Pl(k,n)=\sqcup_{i=0}^{m}\sigma(\widetilde{E}(\lambda^{i})).$
Now $\sigma(\widetilde{E}(\lambda^{i}))$ is equivariantly homeomorphic to $\sigma\mathbb{C}^{d(\lambda^{i})}\times\mathbb{C}^{*}$, where the action of $\mathbb{C}^{*}$ on $\sigma(\widetilde{E}(\lambda^{i}))$ and $\sigma\mathbb{C}^{d(\lambda^{i})}$ are restriction of the ${\bf{b}}$-action on $\mathbb{C}^{m+1}$. The action of $\mathbb{C}^{*}$ on the $\sigma\mathbb{C}^{d(\lambda^{i})}\times\mathbb{C}^{*}$ is given by
$t(z,s)=(tz,t^{b_{\sigma(i)}}s).$
Thus we have the following
$\pi_{b}(\sigma(\widetilde{E}(\lambda^{i})))\cong\frac{\sigma(\widetilde{E}(\lambda^{i}))}{{\bf{b}}\mbox{-action }}\cong\frac{\sigma(\mathbb{C}^{d(\lambda^{i})})}{G(b_{\sigma(i)})}.$
So we get a $q$-CW complex structure of the weighted Grassmann orbifold $\mbox{Gr}_{\bf{b}}(k,n)$ given by
$\mbox{Gr}_{\bf{b}}(k,n)=\frac{\sigma(\mathbb{C}^{d(\lambda^{0})})}{G(b_{\sigma(0)})}\sqcup\frac{\sigma(\mathbb{C}^{d(\lambda^{1})})}{G(b_{\sigma(1)})}\sqcup\dots\sqcup\frac{\sigma(\mathbb{C}^{d(\lambda^{m})})}{G(b_{\sigma(m)})}.$
Also, the usual $q$-CW complex structure on a weighted Grassmann orbifold, as in Subsection 2.2, gives the following $q$-CW complex structure.
$\mbox{Gr}_{\sigma{\bf b}}(k,n)=\frac{\mathbb{C}^{d(\lambda^{0})}}{G(b_{\sigma(0)})}\sqcup\frac{\mathbb{C}^{d(\lambda^{1})}}{G(b_{\sigma(1)})}\sqcup\dots\sqcup\frac{\mathbb{C}^{d(\lambda^{m})}}{G(b_{\sigma(m)})}.$
@@
Let ${\bf{b}}=(b_{0},\dots,b_{m})$ be a Plücker weight vector corresponding to some $(W,a)$ and $\phi\in S_{n}$. Let $W^{\prime}:=(w_{\phi(1)},\dots,w_{\phi(n)})$ and $\{\phi b_{0},\dots,\phi b_{m}\}$ be the Plücker weight corresponding to $(W^{\prime},a)$ as in (2.8). If $\sigma$ is the Plücker permutations on $\{0,1,\dots,m\}$ corresponding to $\phi$ then $\phi b_{i}=b^{\sigma(i)}$. If $\mbox{WGr}^{\prime}(k,n)$ is the weighted Grassmann orbifold corresponding to $(W^{\prime},a)$ then $\mbox{WGr}^{\prime}(k,n)$ is homeomorphic to $\mbox{Gr}_{\sigma{\bf{b}}}(k,n)$. Thus Theorem 3.4 generalizes [7, Theorem 3.5].
Using Definition 2.3 it follows that if ${\bf b}=(b_{0},b_{1},\dots,b_{m})$ is a Plücker weight vector then $r{\bf b}:=(rb_{0},rb_{1},\dots,rb_{m})$ is also a Plücker weight vector for any positive integer $r$.
Proposition 3.6.
@@
Consider the identity map ${\rm id}\colon Pl(k,n)\to Pl(k,n)$. The orbit map
$\pi_{\bf b}\colon Pl(k,n)\rightarrow\mbox{Gr}_{\bf b}(k,n)$
induces an (*$(\mathbb{C}^{*})^{n}$-equivariant homeomorphism $\mbox{Gr}_{\bf b}(k,n)\rightarrow\mbox{Gr}_{r{\bf b}}(k,n)$ for any positive integer $r$ such that the following diagram commutes. The (*$(\mathbb{C}^{*})^{n}$ action is defined in (3.1).
[…]
In [5], the authors have classified the weighted projective spaces $\mathbb{W}P(c_{0},\dots,c_{m})$ up to homeomorphism in terms of weight vectors.
A weight vector $(c_{0},c_{1},\dots,c_{m})\in(\mathbb{Z}_{\geq 0})^{m+1}$ is said to be a primitive weight vector if $\gcd\{c_{0},c_{1},\dots,c_{m}\}=1$. Moreover, A weight vector is called normalized if, for any prime $p$, at least two weights are not divisible by $p$. If a primitive weight vector is not normalized, then a prime $p$ exists such that all weights except one are divisible by $p$. Then, one can apply the normalization process by dividing all those vectors by $p$. One can apply the normalization process recurrently if the new vector is not normalized. Note that a repeated application normalization process can transform any weight vector into a unique normalized vector.

### 09 Mar 2026 v3

@@ 2.3. Lens complex and induced cofibration
(2.15) $I(\lambda^{i})=\{\lambda^{j}:\lambda^{j}=(\lambda^{i}_{s},r)\lambda^{i}\text{ for }(\lambda^{i}_{s},r)\in{\rm{inv}}(\lambda^{i})\text{ and }s=1,2,\dots,k\}.$
Note that $|I(\lambda^{i})|=k(n-k)-d(\lambda^{i})$ for all $i=0,1,\dots,m$. Let $d^{\prime}(\lambda^{i}):=k(n-k)-d(\lambda^{i})$. Consider $\mathbb{C}^{d^{\prime}(\lambda^{i})}\subset\mathbb{C}^{m+1}$ by
(2.16) $\mathbb{C}^{d^{\prime}(\lambda^{i})}=\{(z_{0},z_{1},\dots,z_{m})~|~z_{j}=0\text{ if }\lambda^{j}\notin I(\lambda^{i})\}.$
Then, ${E}(\lambda^{i})$ is homeomorphic to $\mathbb{C}^{d^{\prime}(\lambda^{i})}$ by the following map
$g\colon{E}(\lambda^{i})\to\mathbb{C}^{d^{\prime}(\lambda^{i})}.$
The $j$-th entry of $g([a_{1}\wedge\dots\wedge a_{k}])$ is equal to $a_{sr}$ (the $r$-th entry of $a_{s}$) if $\lambda^{j}=(\lambda^{i}_{s},r)\lambda^{i}$ for some $1\leq s\leq k$, $1\leq r\leq n$ and $r\notin\{\lambda^{i}_{1},\dots,\lambda^{i}_{k}\}$. This defines both $g$ and $g^{-1}$ completely and both $g$ and $g^{-1}$ are continuous.
Consider the $G(b_{i})$-action on $\mathbb{C}^{d^{\prime}(\lambda^{i})}$ by the restriction of the $\bf{b}$}-action of $\mathbb{C}^{*}$ on $\mathbb{C}^{m+1}$ as in (1.1). Recall the $G(b_{i})$ action on $E(\lambda^{i})$ as in (2.12). Then $g$ is a $G(b_{i})$-equivariant homeomorphism. Moreover,
$\frac{E(\lambda^{i})}{G(b_{i})}\cong\frac{\mathbb{C}^{d^{\prime}(\lambda^{i})}}{G(b_{i})}.$
Next, we define weight vectors ${\bf{b}}^{(i)}$ for every $i=0,1,\dots,m$ induced from a Plücker weight vector ${\bf b}=(b_{0},b_{1},\dots,b_{m})$. The weight vector ${\bf{b}}^{(i)}$ is defined by substituting $b_{j}=0$ in $(b_{0},b_{1},\dots,b_{m})$ if $\lambda^{j}\notin I(\lambda^{i})$. Moreover, for every $i=0,1,\dots,m$, we define a lens complex $L^{\prime}(b_{i};{\bf{b}}^{(i)})\subset L^{\prime}(b_{i};(b_{0},\dots,b_{i-1},b_{i+1},\ldots,b_{m}))$ by the following:
$L^{\prime}(b_{i};{\bf{b}}^{(i)}):=\{[(z_{0},z_{1},\dots,z_{m})]\in L^{\prime}(b_{i};(b_{0},\dots,b_{i-1},b_{i+1},\ldots,b_{m}))~|~z_{j}=0\text{ if }\lambda^{j}\notin I(\lambda^{i})\}$
@@
We consider $\mathbb{C}^{d^{\prime}(\lambda^{i})}$ from (2.16) and the unit circle on $\mathbb{C}^{d^{\prime}(\lambda^{i})}$ by
$S^{2d^{\prime}(\lambda^{i})-1}=\{(z_{0},\dots,z_{m})\in\mathbb{C}^{d^{\prime}(\lambda^{i})}~|~\sum_{i=0}^{m}|z_{i}|^{2}=1\}.\text{ Then }$
$L^{\prime}(b_{i};{\bf{b}}^{(i)})=\dfrac{S^{2d^{\prime}(\lambda^{i})-1}}{G(b_{i})}.$
Using the $q$-CW complex decomposition $X^{i}=X^{i-1}\sqcup\frac{\mathbb{C}^{d^{\prime}(\lambda^{i})}}{G(b_{i})}$. Now the result follows from the fact that $\frac{\mathbb{C}^{d^{\prime}(\lambda^{i})}}{G(b_{i})}$ is an unbounded cone over $L^{\prime}(b_{i};{\bf{b}}^{(i)})$. ∎
@@ 3. Classification of weighted Grassmann orbifolds upto certain homeomorphism
In this section, we introduce the Plücker permutation, a permutation on Plücker coordinates up to some sign. Then, we prove that a Plücker permutation induces an equivariant homeomorphism between two weighted Grassmann orbifolds. We also show that two weighted Grassmann orbifolds are homeomorphic if their Plücker weight vectors are differed by a scalar multiplication. We discuss certain equivariant rigidity of weighted Grassmann orbifolds in terms of Plücker weight vector. Moreover, we provide a conjecture that classify weighted Grassmann orbifolds up to homeomorphism.
@@
A permutation $\sigma$ on $\{0,1,\dots,m\}$ is said to be Plücker permutation if there exist a sign vector $t_{\sigma}\in\{1,-1\}^{m+1}$ such that $t_{\sigma}{\cdot}\sigma z\in Pl(k,n)$ for every $z\in Pl(k,n)$.
Let $\sigma$ and $\eta$ be two Plücker permutations on $\{0,1,2,\dots,m\}$. Then $\sigma\circ\eta$ is also a Plücker permutation on $\{0,1,2,\dots,m\}$. Note that, $\sigma^{-1}=\sigma^{i}$ for some positive integer $i$. Thus $\sigma^{-1}$ is also a Plücker permutation on $\{0,1,2,\dots,m\}$. Hence the Plücker permutations on $\{0,1,\dots,m\}$ form a subgroup of the permutation group on $\{0,1,\dots,m\}$.
@@
The permutation group $S_{n}$ can be viewed as a subgroup of the Plücker permutations on $\{0,1,\dots,m\}$ if $m+1={n\choose k}$. To verify this, consider a Schubert symbol $\lambda^{i}=(\lambda_{1}^{i},\ldots,\lambda_{k}^{i})$ for $i\in\{0,1,\ldots,m\}$ and a permutation $\phi\in S_{n}$. Define $\phi\lambda^{i}:=(\phi(\lambda_{i_{1}}^{i}),\ldots,\phi(\lambda_{i_{k}}^{i}))$, where $\{i_{1},\dots,i_{k}\}=\{1,\dots,k\}$ such that $1\leq\phi(\lambda_{i_{1}}^{i})<\cdots<\phi(\lambda_{i_{k}}^{i})\leq n$. Thus $\phi$ gives rise to a permutation $\sigma$ on $\{0,1,\dots,m\}$ defined by $\phi\lambda^{i}=\lambda^{\sigma(i)}$. Moreover, $\sigma$ is a Plücker permutation on $\{0,1,\dots,m\}$ follows from (2.4).
There are many Plücker permutations on $\{0,1,\dots,m\}$ those do no arise from $S_{n}$. For example Consider $d=2$, $n=4$, $m+1=6$. Recall the total order on Schubert symbol as in Example 2.5. Define a permutation $\sigma\in\{0,1,2,3,4,5\}$ by $\sigma(0)=5,\sigma(5)=0$ and $\sigma(i)=i$ for the remaining $i$. Then it is a Plücker permutation corresponding to the sign vector $t_{\sigma}=(1,1,1,1,1,1)$. Note that there does not exist $\phi\in S_{4}$ such that $\phi\lambda^{i}=\lambda^{\sigma(i)}$.
@@
Let ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})$ be a Plücker weight vector. Then $(t^{b_{0}}z_{0},\dots,t^{b_{m}}z_{m})\in Pl(k,n)$ for every $z=(z_{0},\dots,z_{m})\in Pl(k,n)$ and $t\in\mathbb{C}^{*}$. Let $\sigma$ is a Plücker permutation on $\{0,1,\dots,m\}$. Then there exists $t_{\sigma}=(t_{0},\dots,t_{m})\in\{1,-1\}^{m+1}$ such that for every $z=(z_{0},\dots,z_{m})\in Pl(k,n)$ we have $t_{\sigma}{\cdot}\sigma z=(t_{0}z_{\sigma(0)},\dots,t_{m}z_{\sigma(m)})\in Pl(k,n)$ as well as $t_{\sigma}{\cdot}\sigma(t^{b_{0}}z_{0},\dots,t^{b_{m}}z_{m})=(t_{0}t^{b_{\sigma(0)}}z_{\sigma(0)},\dots,t_{m}t^{b_{\sigma(m)}}z_{\sigma(m)})\in Pl(k,n)$. Thus using Remark 2.2, it follows that $\sigma{\bf{b}}=(b_{\sigma(0)},\dots,b_{\sigma(m)})$ is a Plücker weight vector. ∎
Plücker permutations play an important role in characterizing the weighted Grassmann orbifold up to a weakly equivariant homeomorphism. Consider the $(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\bf{b}}(k,n)$ defined by
(3.1) $(t_{1},t_{2},\dots,t_{n})\sum_{i=0}^{m}[a_{i}e_{\lambda^{i}}]=\sum_{i=0}^{m}[t_{\lambda^{i}}a_{i}e_{\lambda^{i}}],$
@@
Lemma 3.4.
If $\sigma$ is a Plücker permutation then $\mbox{Gr}_{\bf{b}}(k,n)$ is weakly equivariantly homeomorphic to $\mbox{Gr}_{\sigma{\bf{b}}}(k,n)$.
@@
The homeomorphism $f_{\sigma}$ becomes weakly equivariant with respect to the permuted $(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\sigma{\bf b}}(k,n)$ defined by
$(t_{1},t_{2},\dots,t_{n})\sum_{i=0}^{m}[a_{i}e_{\lambda^{i}}]=\sum_{i=0}^{m}[t_{\lambda^{\sigma(i)}}a_{i}e_{\lambda^{i}}]$
@@
If $\sigma$ is a Plücker permutation on $\{0,1,\dots,m\}$ then $\Big\{\frac{\mathbb{C}^{d^{\prime}(\lambda^{i})})}{G(b_{\sigma(i)})}\Big\}_{i=0}^{m}$ gives a $q$-CW complex structure of $\mbox{Gr}_{\sigma{\bf b}}(k,n)$ following Subsection $2.$ As a consequence of Lemma 3.4, $\Big\{\frac{\mathbb{C}^{d^{\prime}(\lambda^{i})})}{G(b_{\sigma(i)})}\Big\}_{i=0}^{m}$ gives another $q$-CW complex structure of $\mbox{Gr}_{\bf{b}}(k,n)$.
Example 3.6.
Consider a weighted Grassmann orbifold $\mbox{Gr}_{\bf b}(2,4)$ for the Plücker weight vector ${\bf{b}}=(3,2,1,4,3,2)$ which corresponds to $W=(0,2,1,0)$ and $a=1$. Consider the Plücker permutation $\sigma$ on $\{0,1,2,3,4,5\}$ defined by $\sigma(0)=5,\sigma(5)=0$ and $\sigma(i)=i$ for the remaining $i$. Then $\sigma$ is a Plücker permutation corresponding to $t_{\sigma}=(1,1,1,1,1,1)$. Now $\sigma{\bf{b}}=(2,2,1,4,3,3)$ which corresponds to $W^{\prime}=(-1,1,1,0)$ and $a^{\prime}=2$. Thus both the weighted Grassmann orbifolds $\mbox{WGr}(2,4)$ corresponding to $(W,a)$ and $(W^{\prime},a^{\prime})$ are homeomorphic $\mbox{Gr}_{\bf b}(2,4)$ described above.
Using Definition 2.3, it follows that if ${\bf b}=(b_{0},b_{1},\dots,b_{m})$ is a Plücker weight vector then $r{\bf b}:=(rb_{0},rb_{1},\dots,rb_{m})$ is also a Plücker weight vector for any positive integer $r$.
Lemma 3.7.
@@
Consider the identity map ${\rm id}\colon Pl(k,n)\to Pl(k,n)$. The quotient map
$\pi_{r\bf b}\colon Pl(k,n)\rightarrow\mbox{Gr}_{r\bf b}(k,n)$
induces a map $f$ from $\mbox{Gr}_{r\bf b}(k,n)$ to $\mbox{Gr}_{{\bf b}}(k,n)$ for any positive integer $r$ such that the following diagram commutes.
${Pl(k,n)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rm{id}}$$\scriptstyle{\pi_{r{\bf b}}}$${Pl(k,n)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{{\bf b}}}$${\mbox{Gr}_{r\bf b}(k,n)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}\scriptstyle{\cong}$${\mbox{Gr}_{\bf b}(k,n).}$
Let $U$ be an open subset of $\mbox{Gr}_{\bf b}(k,n)$. Then $(\pi_{\bf b})^{-1}(U)=(\pi_{r{\bf b}})^{-1}\circ f^{-1}(U)$. Since $\pi_{\bf b}$ is a quotient map, the set $(\pi_{\bf b})^{-1}(U)$ is open in $Pl(k,n)$. Thus $f^{-1}(U)$ is an open subset of $\mbox{Gr}_{r{\bf b}}(k,n)$, because $\pi_{r{\bf b}}$ is also a quotient map. Therefore, $f$ is continuous. By a similar argument we can define a map $g$ from $\mbox{Gr}_{\bf b}(k,n)$ to $\mbox{Gr}_{r\bf b}(k,n)$ such that
$f\circ g={\rm id}_{\mbox{Gr}_{\bf b}(k,n)},\text{ and }g\circ f={\rm id}_{\mbox{Gr}_{r\bf b}(k,n)}.$
The map $g=f^{-1}$ is also continuous by the same argument. The $(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\bf b}(k,n)$ and $\mbox{Gr}_{r{\bf b}}(k,n)$ are defined in (3.1). Hence $f$ becomes a $(\mathbb{C}^{*})^{n}$-equivariant homeomorphism. ∎
Theorem 3.8.
Two weighted Grassmann orbifolds $\mbox{Gr}_{\bf b}(k,n)$ and $\mbox{Gr}_{\bf c}(k,n)$ are homeomorphic if the corresponding Plücker weight vectors $\bf b$} and $\bf c$} differ by a Plücker permutation and a Scalar multiplication.
Proof.
The proof follows from Lemma 3.7 together Lemma 3.4. ∎
A weight vector $(c_{0},c_{1},\dots,c_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ is said to be a primitive weight vector if $\gcd\{c_{0},c_{1},\dots,c_{m}\}=1$. Moreover, a primitive weight vector $(c_{0},c_{1},\dots,c_{m})$ is said to be normalized if, for any prime $p$, there exist weights $c_{i}$ and $c_{j}$ for $i\neq j$ such that $p$ does not divides both $c_{i}$ and $c_{j}$. If a primitive weight vector is not normalized, then we can always transform it into a normalized vector which is unique up to an order.